OpenStudy (anonymous):

solve: log(x+5)=log(4x-4)

OpenStudy (ghazi):

use \[\log a - \log b= \log \frac{ a }{ b }\]

OpenStudy (ghazi):

\[\log(x+5)-\log(4x-4)=1\]

OpenStudy (ghazi):

\[\log \frac{ (x+5) }{ (4x-4) }=1\]

OpenStudy (raden):

cancel out the log's both sides : x+5 = 4x - 4 now, solve for x

OpenStudy (anonymous):

Thanks guys! This really helped! :)

OpenStudy (raden):

you're welcome so, what's ur answer ?

OpenStudy (anonymous):

I got x=3

OpenStudy (raden):

yup, u are right